OpenStudy (oswinoswald):

Write the equation in slope-intercept form. What are the slope and y-intercept? -3x - 10y = 7

OpenStudy (oswinoswald):

@mathmath333 @morganchavez @Ashleyisakitty @tester97 @ganeshie8 @undeadknight26

OpenStudy (mathmath333):

y should be on left hand side and everything on right hand side

OpenStudy (mathmath333):

-3x - 10y = 7 - 10y = 7+3x y = -7/10+3x/10

jimthompson5910 (jim_thompson5910):

mathmath333, that +3x/10 should be -3x/10 since you divide both sides by -10

OpenStudy (oswinoswald):

So slope is 3/10 and y-intercept is -7/10?

OpenStudy (mathmath333):

oh yes \[ y = -7/10-3x/10\\ y=-3x/10-7/10\] where m =-3/10 compare it with y=mx+c